Thandi deposits 800 into a bank .The bank will pay simpl interest rate 8% per year.How much will thandi get when she withdraws all her money after 5 years


1
Expert's answer
2022-03-07T19:08:02-0500

"A=P(1+r.n)"

P=800

r=0.08

n=5 years

"=800(1+0.08\u00d75)=1120"